# Branching Brownian motion versus Random Energy Model in the supercritical phase: overlap distribution and temperature susceptibility

### v2

@@ 1.3 Main results
For any $\beta>\beta_{c}=1$, one has $\kappa_{d}(\beta)>\kappa(\beta)$. Moreover, we have
$\frac{3}{2\pi^{2}}-\frac{1}{8}<\liminf_{\beta\downarrow 1}\,(\beta-1)^{2}\,\kappa_{d}(\beta)\leq\limsup_{\beta\downarrow 1}\,(\beta-1)^{2}\,\kappa_{d}(\beta)\leq\frac{3}{\pi^{2}}-\frac{1}{8}.$
As for the two-temperature overlap, we also investigate the general decorated case. Proposition 4.5 shows that $\kappa_{d}(\beta)$ can be written as the sum of $\kappa(\beta)$ and a nonnegative term, which is positive for most non-deterministic decoration processes. Showing that this additional term is growing like $(\beta-1)^{-2}$ as $\beta\downarrow 1$ for the BBM, and in particular the lower bound, is the main task in the proof of Theorem 1.5. In general, this term can be negligible w.r.t. $(\beta-1)^{-2}$ for “small enough” decorations, and it can be much larger than $(\beta-1)^{-2}$ for “big enough” decorations, see Example 4.6. So again, the BBM seems to belong to a critical window, even if here this critical window seems to include more models than for the two-temperature overlap, see e.g. the first family of decorations considered in Example 4.6.
@@ 1.4 Organization of the paper
Section 2 includes preliminary results on the partition function of the REM, on a change of measures used to study the decorated case and on the decoration process of the BBM. Section 3 is dedicated to the two-temperature overlap, including proofs of Theorems 1.1 and 1.2, as well as results in more general decorated cases in Section 3.1. In Section 4, we investigate the temperature susceptibility, proving Theorems 1.3 and 1.5. Appendix A contains the proof of some results needed about the decoration process of the BBM throughout the paper.
@@ 2.1 Properties of $Z(\beta)$
Recall that the partition function of the REM is given by $Z(\beta)=\sum\mathrm{e}^{\beta\xi_{k}}$ where $(\xi_{k})_{k\geq 1}$ are the atoms ranked in non-increasing order of a PPP($\mathrm{e}^{-x}\mathrm{d}x$). It will be convenient to use the notation $\eta_{k}\coloneqq\mathrm{e}^{-\xi_{k}}$ because these are the atoms of a homogeneous Poisson point process on $\mathbb{R}_{+}$ and allows us to use the law of large numbers. Hence $Z(\beta)$, can be rewritten
@@
We first show that $Z(\beta)$ has a stable distribution. We use the following convention: the stable distribution $S_{\alpha}\mathopen{}\mathclose{{}\left(c,s,\mu}\right)$ [^3] Here we do not use $\beta$ for the skewness parameter because the letter is used for the inverse temperature. has a characteristic function given by
@@
In other words, the distribution of $Z(\beta)$ is $S_{1/\beta}\mathopen{}\mathclose{{}\left(\big{(}\Gamma(1-\frac{1}{\beta})\cos\frac{\pi}{2\beta}\big{)}^{\beta},1,0}\right)$.
@@
Lemma 2.2.
@@
$Z(\beta)=\frac{1}{\beta-1}+S+o(1),\quad\text{almost surely,}$,
where $S\sim S_{1}\mathopen{}\mathclose{{}\left(\frac{\pi}{2},1,1-\gamma}\right)$ and $\gamma$ is the Euler constant.
@@ 2.2 Change of measure in the general decorated case
where the $(\xi_{i})_{i\geq 1}$ are the atoms of a PPP($\mathrm{e}^{-x}\mathop{}\mathopen{}\mathrm{d}x$) independent of $(\sum_{k\geq 0}\delta_{d_{ik}})_{i\geq 1}$, which are i.i.d. copies of a point process $\mathcal{D}=\sum_{k\geq 0}\delta_{d_{k}}$ on $(-\infty,0]$ which has a.s. an atom at 0, but we do not assume that $\mathcal{D}$ is the decoration process of the BBM. In particular, for $\beta,\beta^{\prime}>\beta_{c}=1$ the limiting partition function $Z_{d}(\beta)$ is defined as in Equation (1.5) and the limiting mass at 1 of the two-temperature overlap $Q_{d}(\beta,\beta^{\prime})$ is defined as in Equation (1.3).
Moreover, we set
@@
This can be obtained via a direct Laplace transform calculation, or as a consequence of [44, Lemma 2.1] applied with $u_{i}=\mathrm{e}^{\beta\xi_{i}}$, $m=1/\beta$, $X_{i}=S_{\beta,i}$ and $Y_{i}=\sum_{k\geq 0}\delta_{d_{ik}}$. ∎
@@
$Z_{d}(\beta)\coloneqq\sum_{i\geq 1}\sum_{k\geq 0}\mathrm{e}^{\beta(\xi_{i}+d_{ik})}=\sum_{i\geq 1}\mathrm{e}^{\beta\xi_{i}}S_{\beta,i}.$
However, note that one cannot relate the joint distribution of $(Z_{d}(\beta),Z_{d}(\beta^{\prime}))$ with the one of $(Z(\beta),Z(\beta^{\prime}))$ in such a way (otherwise the two-temperature overlap distribution would not depend on the decoration).
Corollary 2.7.
@@
$Z_{d}(\beta)\overset{(\mathrm{d})}{=}\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{1/\beta}}\right]^{\beta}Z(\beta).$
@@
By Lemma 2.6, $Z_{d}(\beta)$ under $\mathbb{P}$ has the same distribution as $\mathrm{e}^{\beta c_{\beta}}Z(\beta)$ under $\mathbb{P}_{\beta}$. But $Z(\beta)$ has the same law under $\mathbb{P}$ and under $\mathbb{P}_{\beta}$, so the result follows. ∎
@@
When $\mathbb{E}[S_{\beta}^{1/\beta}]=\infty$, one can show that $Z_{d}(\beta)=\infty$ a.s. This follows for example from the previous result applied to $S_{\beta}\wedge M$ and then letting $M\to\infty$. Therefore, the assumption that $\mathbb{E}[S_{\beta}^{1/\beta}]<\infty$ for any $\beta>1$ is a very minimalist one when working with the decorated partition function.
@@ 2.3 The decoration of the BBM
appearing in the limit of the extremal process of the BBM. As mentioned in the introduction (see Equation (1.1)), convergence of this extremal process has been established in [1] and [3], which give two different descriptions of the law of $\mathcal{D}$. However, we use here a third description obtained in [21, Lemma 5.1], as well as several other results shown in this paper and its sequel [22]. Note that the authors work with a BBM with branching rate 1 (instead of $1/2$ for us here), but that both processes have the same distribution up to a time-space scaling. In particular, one can check that the decoration process $\mathcal{C}$ appearing in this case (branching rate 1) can be related to $\mathcal{D}$ as follows:
@@
Then, for any positive function $t\mapsto r_{t}$ such that both $r_{t}$ and $t-r_{t}$ tend to $\infty$ as $t\to\infty$, [21, Lemma 5.1] with $u=0$ establishes that, for the vague convergence on the set of Radon measures on $\mathbb{R}$,
@@
Some of the main results in [21] concern the level sets of the decoration point process itself. Recall that $\mathcal{C}$ is supported on $(-\infty,0]$ and has a.s. an atom at 0. In [21, Proposition 1.5], they prove the existence of constants $C_{\star},C>0$ such that
@@
The second moment bound, see Equation (2.9), is actually of the right order, as proved by Cortines, Hartung and Louidor in a second paper [22, Proposition 1.1], which means that the first and second moments of $\mathcal{C}([-x,0])$ are dominated by an unlikely event, as $x\to\infty$. This event already appeared in the proof of Equation (2.8) in [21], when working with $\mathcal{C}_{t,r_{t}}^{*}([-x,0])$ under ~$\widetilde{\mathbb{P}}_{t}$, before taking the limit $t\to\infty$. It consists in the fact that the spine is going sufficiently high at a time $t-s$ with $s$ of order $x^{2}$, more precisely, it can be defined, for some small $\eta>0$ and large $M>0$, as
@@
or in the same way but with $[-M,\infty)$ instead of $[-M,M]$444 To see that both choices are roughly equivalent, note that, if $h_{t-s}(X_{t-s})\simeq m_{t}-m_{s}+y$, then the particles branching from the spine around time $s$ typically (under $\widetilde{\mathbb{P}}$) have descendants at time $t$ close to $m_{t}+y$. Hence, the conditioning in the definition of ~$\widetilde{\mathbb{P}}_{t}$ implies that $h_{t-s}(X_{t-s})$ cannot be much larger than $m_{t}-m_{s}$. . This event has a probability of order $1/x$ and, on this event, $\mathcal{C}_{t,r_{t}}^{*}([-x,0])$ is typically of order $x\mathrm{e}^{\sqrt{2}x}$. Moreover, it is the dominating event in the first and second moments of $\mathcal{C}_{t,r_{t}}^{*}([-x,0])$ and this implies, up to multiplicative constants, that
@@
in accordance with Equations (2.8) and (2.9). See [22, Section 1.3] for a more detailed version of this heuristic picture, which plays a crucial role in the arguments used in this paper for the proofs of Theorems 1.2 and 1.5. We also mention here that the question of the typical size of $\mathcal{C}([-x,0])$ has been investigated in the physics literature [41, 37], where it is conjectured that it differs from $\mathrm{e}^{\sqrt{2}x}$ by a stretched-exponentially small factor $\mathrm{e}^{-\xi x^{2/3}}$ where $\xi$ is a positive random variable. This result has been proved after the first version of this paper in [32].
@@
Recall from Corollary 2.7 that $Z_{d}(\beta)$ has the same distribution as $\mathbb{E}[S_{\beta}^{1/\beta}]^{\beta}Z(\beta)$. On the other hand, we have seen in Lemma 2.4 that $Z(\beta)=\frac{1}{\beta-1}+S+o(1)$ almost surely, as $\beta\downarrow 1$, where $S\sim S_{1}\mathopen{}\mathclose{{}\left(\frac{\pi}{2},1,1-\gamma}\right)$. Combining this with Corollary 2.12 above, we get the following fluctuation result for $Z_{d}(\beta)$:
$(\beta-1)\mathopen{}\mathclose{{}\left(Z_{d}(\beta)-\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{1/\beta}}\right]^{\beta}}{\beta-1}}\right)\xrightarrow[\beta\downarrow 1]{(\mathrm{d})}C_{\star}S.$
Note that the deterministic centering term used here for $Z_{d}(\beta)$ is not explicit: it is asymptotically equivalent to $C_{\star}/(\beta-1)^{2}$ but could involve other non-negligible terms. As mentioned in Remark 2.5, these 1-stable fluctuations appear for other quantities describing the front of the BBM. The content of this remark has been suggested by Xinxin Chen.
@@ 3.1 Results in the general decorated case
We study here the influence of the decorations on the behavior of $F_{d}(\beta,\beta^{\prime})=\mathbb{E}[Q_{d}(\beta,\beta^{\prime})]$, when $\beta^{\prime}>\beta_{c}=1$ is fixed and $\beta\downarrow 1$ and show that they can change the leading order drastically. We work in the general decorated case (see Section 2.2) in order to highlight the fact that the BBM has a critical behavior in some sense to be made precise below.
@@
Recall from Corollary 2.7 and Remark 2.8 that it is a necessary and sufficient condition to have $Z_{d}(\beta)<\infty$ almost surely. In particular, this ensures that $Q_{d}(\beta,\beta^{\prime})$, introduced in Equation (1.3) and which can be written as
$Q_{d}(\beta,\beta^{\prime})=\frac{\sum_{i}\mathrm{e}^{(\beta+\beta^{\prime})\xi_{i}}S_{\beta,i}S_{\beta^{\prime},i}}{\mathopen{}\mathclose{{}\left(\sum_{i}\mathrm{e}^{\beta\xi_{i}}S_{\beta,i}}\right)\mathopen{}\mathclose{{}\left(\sum_{i}\mathrm{e}^{\beta^{\prime}\xi_{i}}S_{\beta^{\prime},i}}\right)},$ (3.2)
is well-defined, and henceforth $F_{d}(\beta,\beta^{\prime})$ as well.
@@
Note that $\mathbb{E}[T_{\beta}]=1$. The fact that the assumptions of the following results are stated in terms of $T_{\beta}$ shows that the behavior of $F_{d}(\beta,\beta^{\prime})$ as $\beta\downarrow 1$ is not governed by the growth rate of $\mathbb{E}[S_{\beta}^{1/\beta}]$ as $\beta\downarrow 1$, but rather by how much $S_{\beta}^{1/\beta}$ is fluctuating around its mean. In the forthcoming assumptions, these fluctuations are controlled in terms of how fast $\mathbb{E}[T_{\beta}^{1+\varepsilon}]$ explodes or $\mathbb{E}[T_{\beta}^{1-\varepsilon}]$ vanishes.
@@
Theorem 3.1.
Assume (3.1) and the following
@@
$F_{d}(\beta,\beta^{\prime})=(1-\alpha+o(1))(\beta-1)\log\frac{1}{\beta-1}.$
@@
$F_{d}(\beta,\beta^{\prime})=(1-\alpha)(\beta-1)\log\frac{1}{\beta-1}+O(\beta-1).$
Corollary 3.2.
@@
$F_{d}(\beta,\beta^{\prime})=(\beta-1)\log\frac{1}{\beta-1}+O(\beta-1).$
@@
In the case $\alpha=1$, which we call the critical case, the previous theorem yields an upper bound for $F_{d}(\beta,\beta^{\prime})$, but does not identify the main order. The following result proves a lower bound.
Theorem 3.3.
@@
$\liminf_{\beta\downarrow 1}\frac{F_{d}(\beta,\beta^{\prime})}{\beta-1}>0.$
Finally, we consider strongly fluctuating cases where $F_{d}(\beta,\beta^{\prime})$ can vanish faster than $(\beta-1)$, as $\beta\downarrow 1$. In that case, the function $\beta\mapsto F_{d}(\beta,\beta^{\prime})$ has a zero derivative at $\beta=1$.
Theorem 3.4.
@@
$F_{d}(\beta,\beta^{\prime})=O\mathopen{}\mathclose{{}\left((\beta-1)^{1+\eta}}\right),\qquad\beta\downarrow 1.$
@@
    $F_{d}(\beta,\beta^{\prime})\sim(1-\alpha)(\beta-1)\log\frac{1}{\beta-1},\qquad\beta\downarrow 1.$
@@
    $F_{d}(\beta,\beta^{\prime})\asymp(\beta-1),\qquad\beta\downarrow 1.$
@@
    $F_{d}(\beta,\beta^{\prime})\leq(\beta-1)^{1+\eta_{0}+o(1)}.$

### v3

@@ 1.3 Main results
For any $\beta>\beta_{c}=1$, one has $\kappa_{\mathrm{d}}(\beta)>\kappa(\beta)$. Moreover, we have
$\frac{3}{2\pi^{2}}-\frac{1}{8}<\liminf_{\beta\downarrow 1}\,(\beta-1)^{2}\,\kappa_{\mathrm{d}}(\beta)\leq\limsup_{\beta\downarrow 1}\,(\beta-1)^{2}\,\kappa_{\mathrm{d}}(\beta)\leq\frac{3}{\pi^{2}}-\frac{1}{8}.$
As for the two-temperature overlap, we also investigate the general decorated case. Proposition 4.5 shows that $\kappa_{\mathrm{d}}(\beta)$ can be written as the sum of $\kappa(\beta)$ and a nonnegative term, which is positive for most non-deterministic decoration processes. Showing that this additional term is growing like $(\beta-1)^{-2}$ as $\beta\downarrow 1$ for the BBM, and in particular the lower bound, is the main task in the proof of Theorem 1.7. In general, this term can be negligible w.r.t. $(\beta-1)^{-2}$ for “small enough” decorations, and it can be much larger than $(\beta-1)^{-2}$ for “big enough” decorations, see Example 4.6. So again, the BBM seems to belong to a critical window, even if here this critical window seems to include more models than for the two-temperature overlap, see e.g. the first family of decorations considered in Example 4.6.
Finally we state two open questions similar to the ones for the two-temperature overlap.
Open question 1.8 (BBM case).
Does the limit of $(\beta-1)^{2}\,\kappa_{\mathrm{d}}(\beta)$ as $\beta\downarrow 1$ exists?
In the context of susceptibility, the number of levels of replica symmetry breaking does not really matter. However, one can expect that the susceptibility is also explicit for the GREM and it would be interesting to compare it with the one for the REM.
Open question 1.9.
Find the susceptibility for the GREM.
@@ 1.4 Organization of the paper
Section 2 includes preliminary results on the partition function of the REM, on a change of measures used to study the decorated case and on the decoration process of the BBM. Section 3 is dedicated to the two-temperature overlap, including proofs of Theorems 1.1 and 1.2, as well as results in more general decorated cases in Section 3.1. In Section 4, we investigate the temperature susceptibility, proving Theorems 1.5 and 1.7. Appendix A contains the proof of some results needed about the decoration process of the BBM throughout the paper.
@@ 2.1 Properties of $Z(\beta)$
In this section, we collect some basic facts on the partition function of the REM in the supercritical phase $\beta>1$ that will be useful in the following sections. We work directly with the limiting extremal process in the spirit of Ruelle’s reformulation of the REM [51] (see also [31] by Derrida and Mottishaw where this model is called the Poisson REM) : the partition function is given by $Z(\beta)=\sum\mathrm{e}^{\beta\xi_{k}}$ where $(\xi_{k})_{k\geq 1}$ are the atoms ranked in non-increasing order of a PPP($\mathrm{e}^{-x}\mathrm{d}x$). The fact that this is the limit of the properly renormalized partition function of the finite size REM is proved by Bovier, Kurkova and Löwe [18, Theorem 1.5].
It will be convenient to use the notation $\eta_{k}\coloneqq\mathrm{e}^{-\xi_{k}}$ because these are the atoms of a homogeneous Poisson point process on $\mathbb{R}_{+}$ and allows us to use the law of large numbers. Hence $Z(\beta)$, can be rewritten
@@
We first show that $Z(\beta)$ has a stable distribution. We use the following convention: the stable distribution ${\rm Stable}_{\alpha}\mathopen{}\mathclose{{}\left(c,s,\mu}\right)$ [^3] Here we do not use $\beta$ for the skewness parameter because the letter is used for the inverse temperature. has a characteristic function given by
@@
In other words, the distribution of $Z(\beta)$ is ${\rm Stable}_{1/\beta}\mathopen{}\mathclose{{}\left(\big{(}\Gamma(1-\frac{1}{\beta})\cos\frac{\pi}{2\beta}\big{)}^{\beta},1,0}\right)$.
@@
Remark 2.2.
The fact that the partition function has a stable distribution appears as a consequence of [6, Theorem 2.3] by Ben Arous, Bogachev and Molchanov, where the authors study the limiting distributions of exponential sums of i.i.d. random variables.
Lemma 2.3 (REM case).
@@
$Z(\beta)=\frac{1}{\beta-1}+W+o(1),\quad\text{almost surely,}$,
where $W\sim{\rm Stable}_{1}\mathopen{}\mathclose{{}\left(\frac{\pi}{2},1,1-\gamma}\right)$ and $\gamma$ is the Euler constant.
@@ 2.2 Change of measure in the general decorated case
where the $(\xi_{i})_{i\geq 1}$ are the atoms of a PPP($\mathrm{e}^{-x}\mathop{}\mathopen{}\mathrm{d}x$) independent of $(\sum_{k\geq 0}\delta_{d_{ik}})_{i\geq 1}$, which are i.i.d. copies of a point process $\mathcal{D}=\sum_{k\geq 0}\delta_{d_{k}}$ on $(-\infty,0]$ which has a.s. an atom at 0, but we do not assume that $\mathcal{D}$ is the decoration process of the BBM. In particular, for $\beta,\beta^{\prime}>\beta_{c}=1$ the limiting partition function $Z_{\mathrm{d}}(\beta)$ is defined as in Equation (1.5) and the limiting mass at 1 of the two-temperature overlap $Q_{\mathrm{d}}(\beta,\beta^{\prime})$ is defined as in Equation (1.3).
Moreover, let us introduce the partition functions associated to the decoration processes [^4] Note that we do not add a subscript “d” in this notation for the sake of readability in the proofs where the quantity $S_{\beta}$ is ubiquitous. There is no confusion possible with the REM for which the quantity $S_{\beta}$ does not appear (it is equal to 1 in that case).:
@@
This can be obtained via a direct Laplace transform calculation, or as a consequence of [49, Lemma 2.1] applied with $u_{i}=\mathrm{e}^{\beta\xi_{i}}$, $m=1/\beta$, $X_{i}=S_{\beta,i}$ and $Y_{i}=\sum_{k\geq 0}\delta_{d_{ik}}$. ∎
@@
$Z_{\mathrm{d}}(\beta)\coloneqq\sum_{i\geq 1}\sum_{k\geq 0}\mathrm{e}^{\beta(\xi_{i}+d_{ik})}=\sum_{i\geq 1}\mathrm{e}^{\beta\xi_{i}}S_{\beta,i}.$
However, note that one cannot relate the joint distribution of $(Z_{\mathrm{d}}(\beta),Z_{\mathrm{d}}(\beta^{\prime}))$ with the one of $(Z(\beta),Z(\beta^{\prime}))$ in such a way (otherwise the two-temperature overlap distribution would not depend on the decoration).
Corollary 2.8 (General decorated case).
@@
$Z_{\mathrm{d}}(\beta)\overset{(\mathrm{d})}{=}\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{1/\beta}}\right]^{\beta}Z(\beta).$
@@
By Lemma 2.7, $Z_{\mathrm{d}}(\beta)$ under $\mathbb{P}$ has the same distribution as $\mathrm{e}^{\beta c_{\beta}}Z(\beta)$ under $\mathbb{P}_{\beta}$. But $Z(\beta)$ has the same law under $\mathbb{P}$ and under $\mathbb{P}_{\beta}$, so the result follows. ∎
@@
When $\mathbb{E}[S_{\beta}^{1/\beta}]=\infty$, one can show that $Z_{\mathrm{d}}(\beta)=\infty$ a.s. This follows for example from the previous result applied to $S_{\beta}\wedge M$ and then letting $M\to\infty$. Therefore, the assumption that $\mathbb{E}[S_{\beta}^{1/\beta}]<\infty$ for any $\beta>1$ is a very minimalist one when working with the decorated partition function.
@@ 2.3 The decoration of the BBM
appearing in the limit of the extremal process of the BBM. As mentioned in the introduction (see Equation (1.1)), convergence of this extremal process has been established in [1] and [3], which give two different descriptions of the law of $\mathcal{D}$. However, we use here a third description obtained in [26, Lemma 5.1], as well as several other results shown in this paper and its sequel [27]. Note that the authors work with a BBM with branching rate 1 (instead of $1/2$ for us here), but that both processes have the same distribution up to a time-space scaling. In particular, one can check that the decoration process $\mathcal{C}$ appearing in this case (branching rate 1) can be related to $\mathcal{D}$ as follows:
@@
Then, for any positive function $t\mapsto r_{t}$ such that both $r_{t}$ and $t-r_{t}$ tend to $\infty$ as $t\to\infty$, [26, Lemma 5.1] with $u=0$ establishes that, for the vague convergence on the set of Radon measures on $\mathbb{R}$,
@@
Some of the main results in [26] concern the level sets of the decoration point process itself. Recall that $\mathcal{C}$ is supported on $(-\infty,0]$ and has a.s. an atom at 0. In [26, Proposition 1.5], they prove the existence of constants $C_{\star},C>0$ such that
@@
The second moment bound, see Equation (2.9), is actually of the right order, as proved by Cortines, Hartung and Louidor in a second paper [27, Proposition 1.1], which means that the first and second moments of $\mathcal{C}([-x,0])$ are dominated by an unlikely event, as $x\to\infty$. This event already appeared in the proof of Equation (2.8) in [26], when working with $\mathcal{C}_{t,r_{t}}^{*}([-x,0])$ under ~$\widetilde{\mathbb{P}}_{t}$, before taking the limit $t\to\infty$. It consists in the fact that the spine is going sufficiently high at a time $t-s$ with $s$ of order $x^{2}$, more precisely, it can be defined, for some small $\eta>0$ and large $M>0$, as
@@
or in the same way but with $[-M,\infty)$ instead of $[-M,M]$555 To see that both choices are roughly equivalent, note that, if $h_{t-s}(X_{t-s})\simeq m_{t}-m_{s}+y$, then the particles branching from the spine around time $s$ typically (under $\widetilde{\mathbb{P}}$) have descendants at time $t$ close to $m_{t}+y$. Hence, the conditioning in the definition of ~$\widetilde{\mathbb{P}}_{t}$ implies that $h_{t-s}(X_{t-s})$ cannot be much larger than $m_{t}-m_{s}$. . This event has a probability of order $1/x$ and, on this event, $\mathcal{C}_{t,r_{t}}^{*}([-x,0])$ is typically of order $x\mathrm{e}^{\sqrt{2}x}$. Moreover, it is the dominating event in the first and second moments of $\mathcal{C}_{t,r_{t}}^{*}([-x,0])$ and this implies, up to multiplicative constants, that
@@
in accordance with Equations (2.8) and (2.9). See [27, Section 1.3] for a more detailed version of this heuristic picture, which plays a crucial role in the arguments used in this paper for the proofs of Theorems 1.2 and 1.7. We also mention here that the question of the typical size of $\mathcal{C}([-x,0])$ has been investigated in the physics literature [46, 42], where it is conjectured that it differs from $\mathrm{e}^{\sqrt{2}x}$ by a stretched-exponentially small factor $\mathrm{e}^{-\xi x^{2/3}}$ where $\xi$ is a positive random variable. This result has been proved after the first version of this paper in [37].
@@
Recall from Corollary 2.8 that $Z_{\mathrm{d}}(\beta)$ has the same distribution as $\mathbb{E}[S_{\beta}^{1/\beta}]^{\beta}Z(\beta)$. On the other hand, we have seen in Lemma 2.5 that $Z(\beta)=\frac{1}{\beta-1}+W+o(1)$ almost surely, as $\beta\downarrow 1$, where $W\sim{\rm Stable}_{1}\mathopen{}\mathclose{{}\left(\frac{\pi}{2},1,1-\gamma}\right)$. Combining this with Corollary 2.13 above, we get the following fluctuation result for $Z_{\mathrm{d}}(\beta)$:
$(\beta-1)\mathopen{}\mathclose{{}\left(Z_{\mathrm{d}}(\beta)-\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{1/\beta}}\right]^{\beta}}{\beta-1}}\right)\xrightarrow[\beta\downarrow 1]{(\mathrm{d})}C_{\star}\,W.$
Note that the deterministic centering term used here for $Z_{\mathrm{d}}(\beta)$ is not explicit: it is asymptotically equivalent to $C_{\star}/(\beta-1)^{2}$ but could involve other non-negligible terms. As mentioned in Remark 2.6, these 1-stable fluctuations appear for other quantities describing the front of the BBM. The content of this remark has been suggested by Xinxin Chen.
@@ 3.1 Results in the general decorated case
We study here the influence of the decorations on the behavior of $\mathbb{E}[Q_{\mathrm{d}}(\beta,\beta^{\prime})]$, when $\beta^{\prime}>\beta_{c}=1$ is fixed and $\beta\downarrow 1$ and show that they can change the leading order drastically. We work in the general decorated case (see Section 2.2) in order to highlight the fact that the BBM has a critical behavior in some sense to be made precise below.
@@
Recall from Corollary 2.8 and Remark 2.9 that it is a necessary and sufficient condition to have $Z_{\mathrm{d}}(\beta)<\infty$ almost surely. In particular, this ensures that $Q_{\mathrm{d}}(\beta,\beta^{\prime})$, introduced in Equation (1.3) and which can be written as
$Q_{\mathrm{d}}(\beta,\beta^{\prime})=\frac{\sum_{i}\mathrm{e}^{(\beta+\beta^{\prime})\xi_{i}}S_{\beta,i}S_{\beta^{\prime},i}}{\mathopen{}\mathclose{{}\left(\sum_{i}\mathrm{e}^{\beta\xi_{i}}S_{\beta,i}}\right)\mathopen{}\mathclose{{}\left(\sum_{i}\mathrm{e}^{\beta^{\prime}\xi_{i}}S_{\beta^{\prime},i}}\right)},$ (3.2)
is well-defined, and henceforth $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ as well.
@@
Note that $\mathbb{E}[T_{\beta}]=1$. The fact that the assumptions of the following results are stated in terms of $T_{\beta}$ shows that the behavior of $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ as $\beta\downarrow 1$ is not governed by the growth rate of $\mathbb{E}[S_{\beta}^{1/\beta}]$ as $\beta\downarrow 1$, but rather by how much $S_{\beta}^{1/\beta}$ is fluctuating around its mean. In the forthcoming assumptions, these fluctuations are controlled in terms of how fast $\mathbb{E}[T_{\beta}^{1+\varepsilon}]$ explodes or $\mathbb{E}[T_{\beta}^{1-\varepsilon}]$ vanishes.
@@
Theorem 3.1 (General decorated case).
Assume (3.1) and that the decoration process $\mathcal{D}$ satisfies the following properties:
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]=(1-\alpha+o(1))(\beta-1)\log\frac{1}{\beta-1}.$
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]=(1-\alpha)(\beta-1)\log\frac{1}{\beta-1}+O(\beta-1).$
Corollary 3.2 (General decorated case).
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]=(\beta-1)\log\frac{1}{\beta-1}+O(\beta-1).$
@@
In the case $\alpha=1$, which we call the critical case, the previous theorem yields an upper bound for $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$, but does not identify the main order. The following result proves a lower bound.
Theorem 3.3 (General decorated case).
@@
$\liminf_{\beta\downarrow 1}\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]}{\beta-1}>0.$
Finally, we consider strongly fluctuating cases where $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ can vanish faster than $(\beta-1)$, as $\beta\downarrow 1$. In that case, the function $\beta\mapsto\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ has a zero derivative at $\beta=1$.
Theorem 3.4 (General decorated case).
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]=O\mathopen{}\mathclose{{}\left((\beta-1)^{1+\eta}}\right),\qquad\beta\downarrow 1.$
@@
    $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]\sim(1-\alpha)(\beta-1)\log\frac{1}{\beta-1},\qquad\beta\downarrow 1.$
@@
    $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]\asymp(\beta-1),\qquad\beta\downarrow 1.$
@@
    $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]\leq(\beta-1)^{1+\eta_{0}+o(1)}.$